If the rate of disappearance of H2 is 3.8 M/s, what is the rate of disappearance of Fe3O4? Answer to 2 decimal places.
Answer #1

The chemical reaction between Fe3O4 and H2 is-

Fe3O4 + 4H2 ----------> 3Fe + 4H2O

i.e 1 mole of Fe3O4 will need to react with 4 moles of H2

Or we can say when 4 moles of H2 disapperas from the reactant, only 1 mole of Fe has been disappeared from it.

SO rate of disappearance of H2 i.e the moles of H2 disapperead per unit time will be 4 times the rate of disappearance of Fe3O4 i.e the moles of Fe3O4 disapperead per unit time

Or  rate of disappearance of Fe3O4 unit time will be 1/4 times the rate of disappearance of H2

So we can write-

rate of disappearance of Fe3O4 = 1/4 * rate of disappearance of H2

- d[Fe3O4] / dt = -1/4 * d[H2] / dt

=  1/4 * 3.8 M/s

= 0.95‬ M/s

So rate of disappearance of Fe3O4 is 0.95‬ M/s